OpenStudy (anonymous):

3. The measures of two of the sides of an equilateral triangle are 3x + 15 in. and 7x – 5 in. What is the measure of the third side in inches? (1 point) 5 1 18 30

OpenStudy (aripotta):

3x + 15 = 7x - 5 15 = 4x - 5 20 = 4x x = 5

OpenStudy (aripotta):

3(5) + 15 = ?

OpenStudy (anonymous):

30?

OpenStudy (aripotta):

yes

OpenStudy (anonymous):

so thats my answer?:)

OpenStudy (aripotta):

yes

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

since its an equilateral triangle, and all the sides of an equilateral triangle are equal, then you equate the two given sides to each other to find for x and substitute it in any of the given equations just as my friend has done